OpenStudy (anonymous):

Find dy/dx in terms of x and y if xlny+y5=7lnx

OpenStudy (anonymous):

implicit diff yes?

OpenStudy (anonymous):

@satellite73 Do You Think You Can Help With A Problem?

OpenStudy (anonymous):

\[ x\ln(y)+y^5=7\ln(x)\] the right side is easy, it is \[\frac{7}{x}\] the left side requires the product and chain rule

OpenStudy (anonymous):

yep yep

OpenStudy (anonymous):

\[\ln(y)+x\frac{1}{y}y'+5y^4y'=\frac{7}{x}\] is a start

OpenStudy (anonymous):

Thanks! I just need a start and can go from there. The online website we have to enter isn't made very well so it's always a struggle entering the answers.

OpenStudy (anonymous):

is it clear how i got that? that is really the only concept part, the rest is algebra

OpenStudy (anonymous):

Yeah it makes sense now. The algebra of it is pretty easy, to start it off is what's confusing

OpenStudy (anonymous):

ok good